Find the critical numbers of the function

ƒ(x) = 2x³ +8x +x²


Find the absolute maximum and minimum values of f on the given interval

ƒ(x) = x^4 + 4x³ - 9



Verify the MVT and find the value of c in the given interval

 y = x³ - 9x² + 24x − 18; [2, 4]


1
Expert's answer
2021-08-19T07:43:31-0400

1.


"f(x)=2x^3+8x+x^2"

The critical numbers of the function are such numbers that "f'(x)=0."

Find "f'(x)"


"f'(x)=6x^2+2x+8"

Solve for x


"6x^2+2x+8=0\newline
3x^2+x+4=0"

"D=1-4\u20223\u20224=-47, D<0,"

there are not real roots.

Answer. There aren't critical numbers.

2.


"f(x)=x^4+4x^3-9"

Find


"f'(x)=4x^3+12x^2"

Solve for x


"4x^3+12x^2=0""4x^2(x+3)=0""x=0, \text{or } x=-3"

There are two critical numbers 0 and -3.

Find the sign of derivative on the intervals "(-\infty,-3),(-3,0),(0,\infty)."

On the interval "(-\infty,-3)" "f'(x)<0, f(x)" decreases.

On the interval "(-3,0)" "f'(x)>0, f(x)" increases.

On the interval "(0,\infty)" "f'(x)>0, f(x)" increases.

Consider the segment [a,b], where a<-3 and b>0. Then the critical point -3 and 0 are in segment [a,b].

Let be a=-5 and b=1

Find

"f(-5)=(-5)^4+4\cdot(-5)^3-9=491
\newline f(0)=-9
\newline f(1)=1^4+4\cdot1^3-9=-4"

"f(-3)=(-3)^4+4\u2022(-3)^3-9=-36."

Answer. Therefore function "f(x)" has the absolute minimum value -36 and the absolute maximum value 491 on the segment [-5,1].


3.

The function f is continuous on the interval [2,4] and differentiable on the interval ]2,4[, therefore we can apply the MVT (mean value theorem) which states that there is c∈[2,4]

 such that


"f'(c)=\frac{f(4)-f(2)}{4-2}""f(x)=x^3-9x^2+24x-18""f(2)=2, f(4)=-2"


Therefore,


"f'(c)=(-2-2)\/2=-2"


Also,


"f'(x)=3x^2-18x+24""f'(c)=3c^2-18c+24=-2"

Solving for c


"3c^2-18c+26=0"

"D=12, D>0," so there are 2 real roots


"c=\frac{9\pm\sqrt{3}}{3}."

Both values of c ∈[2,4].

Answer. "c=\frac{9\pm\sqrt{3}}{3}"